# Asymptotic constraints for 1D planar grey photon diffusion from linear transport with special-relativistic effects

### v1

@@ 4.1 Constant $\beta$
Here we compare the drift-diffusion equation with lab frame MC transport at several constant velocities. This removes the effect of the adiabatic and Doppler shift term in the drift diffusion equation, isolating the effect of the $\gamma$-factors. Figure 2 has scaled comoving scalar flux versus position, where each MC (solid lines) result has its peak normalized to 1, and the corresponding diffusion results are scaled by the same resulting factor, per $\beta$ value. The result for $\beta=0$ merely confirms the static solutions are consistent. At $\beta=0.3$, we see the effect of beaming in the MC and the effect of advection in diffusion cause the solutions to skew to the right. Moreover, at $\beta=0.3$, all three results agree closely. However, at $\beta=0.6$ and 0.9, we see that the semi-relativistic result suffers in comparison to the MC result, consistent with O($\beta^{2}$) effects from the $\gamma$-factors becoming important. The fully relativistic drift-diffusion result agrees closely at each $\beta$, but is missing a spike near the source present in the MC result.
Figure 2: Scaled scalar flux versus spatial coordinate for several values of constant $\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 4.2 Homologous flow, $\beta=\beta_{c}+(x-1/2)\Delta\beta$
Figure 3: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.6+(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 4.3 Velocity jump, $\beta=\beta_{l}+\Theta(x-1/2)\Delta\beta$
Finally, we consider a problem with a very sharp, localized velocity gradient, where our velocity discretization approximates $\beta=0.3+\Theta(x-1/2)\Delta\beta$, where $\Theta(\cdot)$ is again the unit step function. Departing from the previous problems we consider a uniform comoving volume source. The jump in velocity at $x=1/2$ incurs a significant change in the comoving scalar intensity, as seen in Fig. 4 for all three solutions. In Fig. 4 we have MC transport (solid lines), semi-relativistic (dashed lines) and fully-relativistic (dotted lines) diffusion. We see that all three solutions are comparable for the $\Delta\beta=0.3$ jump (to $\beta=0.6$ at $x\geq 1/2$). For the larger jump $\Delta\beta=0.6$, we see that the fully relativistic solution is in better agreement in the higher-velocity region, but all three solutions agree fairly well in the lower-velocity region.
Figure 4: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.3+\Theta(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 5 Conclusions
We have derived a fully-relativistic grey diffusion equation for continuous-direction 1D planar linear radiative transfer using a lab-frame-based asymptotic analysis. This analysis makes use of a family of parameter scaling that is intended to be in proximity to the standard static-material background scaling used in linear transport theory. The resulting fully-relativistic diffusion equation has a diffusion operator that scales as $1/\gamma^{3}$, consistent with the two-direction Poisson-Kac-based derivation of [giona2017].
@@
The standard static-material scaling rules for thermal radiative transfer are somewhat different than linear transport: the absorption opacity is scaled to be asymptotically large and one considers a matter equation with an asymptotically small heat capacity [larsen1983]. The time derivative is scaled as in the linear transport parameter scaling. This approach is evidently distinct from the opacity-only scaling method of [thomas1930, anderson1972] considered in the general-relativistic framework. General relativistic codes often use a time increment that does not correspond to a simple lab-frame picture, however, and the time variable is on more equal footing with the spatial variables (see Appendix Section C).
For 3D, we must consider the finding of [giona2017] that shows an anisotropic effective diffusion coefficient. While this result is still for discrete directions, we anticipate a similar effect for continuous direction in 3D. One possible path forward on a 3D derivation is to consider the analysis on a piecewise constant velocity field, where one can rotate the spatial coordinate system to have an axis aligned with the velocity in each cell; then it may be possible to follow much of the analysis given here, but with an additional step of inverting the spatial rotation (and projecting it onto a stencil that is compatible with the cells).
@@ B.3.3 Neglecting terms $\sim\partial(\beta F)/\partial x$
Equation (116) is the 1D planar form of Eq. 6.51 of [castor2004], but with the following modifications:
@@
However, Eq. 6.51 of [castor2004] is for comoving $E$ and $F$. Similarly, this is the limit we obtain of (48) ((49)) in Section 2.4, but with lab-frame $E$.
@@ Appendix C The Thomas scaling in non-relativistic 1D planar geometry
Here we briefly consider the Thomas scaling [thomas1930], as presented by [anderson1972], in a non-relativistic 1D context with constant pure absorption opacity. To the best of our understanding, the only small parameter is the mean-free path in these approaches. This should constitute a simplification of Section IV of [anderson1972]. Using similar notation to [anderson1972], the transport equation is then
@@
where we have written the partial derivatives of the streaming operator in terms of the characteristic, or affine coordinate $\lambda$. Following [anderson1972], we write the asymptotic expansion directly in powers of $\sigma$.
@@
We note that this is equivalent to Eq. (25), by introducing a length scale $L$, and setting $\varepsilon=1/L\sigma$ and $\psi^{(k)}=L^{k}\tilde{\psi}^{(k)}$. Again following [anderson1972], matching orders to O($1/\sigma$) gives
@@
which is nearly the Telegrapher’s equation (it is so with an imaginary relaxation coefficient of the 2nd derivative in time). Putting into the 4-vector notation of [anderson1972, shibata2011], Eq. (123) becomes
@@
where $\alpha$ and $\delta$ here are 4-indexes, $u^{\alpha}=(1,0,0,0)$ is the 4-velocity, and $h^{\alpha\delta}=u^{\alpha}u^{\delta}-\eta^{\alpha\delta}$ is the rank-4 projection operator, as in [anderson1972, shibata2011] ($\eta^{\alpha\delta}$ here is the Minkowski metric).
In contrast to equation 5.21 of [shibata2011], Eq. (124) has an additional time-like component in the 2nd derivative, which can be seen as a contribution of $\tilde{\phi}^{(1)}$ (consistent with equation 5.13 of [shibata2011]). Excising this time-like component would furnish a non-equilibrium diffusion equation from Eqs. (123) and (124).

### v2

@@ 4.1 Constant $\beta$
Here we compare the drift-diffusion equation with lab frame MC transport at several constant velocities. This problem has a discretized point source $Q_{0,i}=1/\Delta x$ at $x=0.5$. This removes the effect of the adiabatic and Doppler shift term in the drift diffusion equation, isolating the effect of the $\gamma$-factors. Figure 2 has scaled comoving scalar flux versus position, where each MC (solid lines) result has its peak normalized to 1, and the corresponding diffusion results are scaled by the same resulting factor, per $\beta$ value. The result for $\beta=0$ merely confirms the static solutions are consistent. At $\beta=0.3$, we see the effect of beaming in the MC and the effect of advection in diffusion cause the solutions to skew to the right. Moreover, at $\beta=0.3$, all three results agree closely. However, at $\beta=0.6$ and 0.9, we see that the semi-relativistic result suffers in comparison to the MC result, consistent with O($\beta^{2}$) effects from the $\gamma$-factors becoming important. The fully relativistic drift-diffusion result agrees closely at each $\beta$, but is missing a spike near the source present in the MC result.
Figure 2: Scaled scalar flux versus spatial coordinate for a point source at $x=0.5$ and several values of constant $\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 4.2 Homologous flow, $\beta=\beta_{c}+(x-1/2)\Delta\beta$
Figure 3: Scaled scalar flux versus spatial coordinate for a point source at $x=0.5$ and several values of $\Delta\beta=0.06,0.6$ in $\beta=0.6+(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 4.3 Velocity jump, $\beta=\beta_{l}+\Theta(x-1/2)\Delta\beta$
Finally, we consider a problem with a very sharp, localized velocity gradient, where our velocity discretization approximates $\beta=0.3+\Theta(x-1/2)\Delta\beta$, where $\Theta(\cdot)$ is again the unit step function. Departing from the previous problems we consider a uniform comoving volume source, $Q_{0}=1$. The jump in velocity at $x=1/2$ incurs a significant change in the comoving scalar intensity, as seen in Fig. 4 for all three solutions. In Fig. 4 we have MC transport (solid lines), semi-relativistic (dashed lines) and fully-relativistic (dotted lines) diffusion. We see that all three solutions are comparable for the $\Delta\beta=0.3$ jump (to $\beta=0.6$ at $x\geq 1/2$). For the larger jump $\Delta\beta=0.6$, we see that the fully relativistic solution is in better agreement in the higher-velocity region, but all three solutions agree fairly well in the lower-velocity region.
Figure 4: Scaled scalar flux versus spatial coordinate for a uniform comoving source and several values of $\Delta\beta=0.06,0.6$ in $\beta=0.3+\Theta(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@ 5 Conclusions
We have derived a fully-relativistic grey diffusion equation for continuous-direction 1D planar linear radiative transfer using a lab-frame-based asymptotic analysis. This analysis makes use of a family of parameter scaling that is intended to be in proximity to the standard static-material background scaling used in linear transport theory. The resulting fully-relativistic diffusion equation has a diffusion operator that scales as $1/\gamma^{3}$, consistent with the two-direction Poisson-Kac-based derivation of [14].
@@
The standard static-material scaling rules for thermal radiative transfer are somewhat different than linear transport: the absorption opacity is scaled to be asymptotically large and one considers a matter equation with an asymptotically small heat capacity [23]. The time derivative is scaled as in the linear transport parameter scaling. This approach is evidently distinct from the opacity-only scaling method of [34, 3] considered in the general-relativistic framework. General relativistic codes often use a time increment that does not correspond to a simple lab-frame picture, however, and the time variable is on more equal footing with the spatial variables (see Appendix Section C).
For 3D, we must consider the finding of [14] that shows an anisotropic effective diffusion coefficient. While this result is still for discrete directions, we anticipate a similar effect for continuous direction in 3D. One possible path forward on a 3D derivation is to consider the analysis on a piecewise constant velocity field, where one can rotate the spatial coordinate system to have an axis aligned with the velocity in each cell; then it may be possible to follow much of the analysis given here, but with an additional step of inverting the spatial rotation (and projecting it onto a stencil that is compatible with the cells).
@@ B.3.3 Neglecting terms $\sim\partial(\beta F)/\partial x$
Equation (116) is the 1D planar form of Eq. 6.51 of [5], but with the following modifications:
@@
However, Eq. 6.51 of [5] is for comoving $E$ and $F$. Similarly, this is the limit we obtain of (48) ((49)) in Section 2.4, but with lab-frame $E$.
@@ Appendix C The Thomas scaling in non-relativistic 1D planar geometry
Here we briefly consider the Thomas scaling [34], as presented by [3], in a non-relativistic 1D context with constant pure absorption opacity. To the best of our understanding, the only small parameter is the mean-free path in these approaches. This should constitute a simplification of Section IV of [3]. Using similar notation to [3], the transport equation is then
@@
where we have written the partial derivatives of the streaming operator in terms of the characteristic, or affine coordinate $\lambda$. Following [3], we write the asymptotic expansion directly in powers of $\sigma$.
@@
We note that this is equivalent to Eq. (25), by introducing a length scale $L$, and setting $\varepsilon=1/L\sigma$ and $\psi^{(k)}=L^{k}\tilde{\psi}^{(k)}$. Again following [3], matching orders to O($1/\sigma$) gives
@@
which is nearly the Telegrapher’s equation (it is so with an imaginary relaxation coefficient of the 2nd derivative in time). Putting into the 4-vector notation of [3, 31], Eq. (123) becomes
@@
where $\alpha$ and $\delta$ here are 4-indexes, $u^{\alpha}=(1,0,0,0)$ is the 4-velocity, and $h^{\alpha\delta}=u^{\alpha}u^{\delta}-\eta^{\alpha\delta}$ is the rank-4 projection operator, as in [3, 31] ($\eta^{\alpha\delta}$ here is the Minkowski metric).
In contrast to equation 5.21 of [31], Eq. (124) has an additional time-like component in the 2nd derivative, which can be seen as a contribution of $\tilde{\phi}^{(1)}$ (consistent with equation 5.13 of [31]). Excising this time-like component would furnish a non-equilibrium diffusion equation from Eqs. (123) and (124).